OpenStudy (anonymous):

help please finding exact solutions of cos2x+10cosx=5

OpenStudy (ybarrap):

Is this the problem? $$ \cos(2x)+10\cos x=5 $$

OpenStudy (anonymous):

yes and you have to find solutions

OpenStudy (anonymous):

i tryed to use the identity for cos(2x) but it didnt work

OpenStudy (ybarrap):

There are real and imaginary solutions: http://www.wolframalpha.com/input/?i=cos%282x%29%2B10cosx%3D5 Here are the details for the real solution: Solve for x: 10 cos(x)+cos(2 x) = 5 Simplify the left hand side. Transform 10 cos(x)+cos(2 x) into a polynomial with respect to cos(x) using cos(2 x) = 2 cos^2(x)-1: -1+10 cos(x)+2 cos^2(x) = 5 Write the quadratic equation in standard form. Divide both sides by 2: -1/2+5 cos(x)+cos^2(x) = 5/2 Solve the quadratic equation by completing the square. Add 1/2 to both sides: 5 cos(x)+cos^2(x) = 3 Take one half of the coefficient of cos(x) and square it, then add it to both sides. Add 25/4 to both sides: 25/4+5 cos(x)+cos^2(x) = 37/4 Factor the left hand side. Write the left hand side as a square: (5/2+cos(x))^2 = 37/4 Eliminate the exponent on the left hand side. Take the square root of both sides: 5/2+cos(x) = sqrt(37)/2 or 5/2+cos(x) = -sqrt(37)/2 Look at the first equation: Solve for cos(x). Subtract 5/2 from both sides: cos(x) = sqrt(37)/2-5/2 or 5/2+cos(x) = -sqrt(37)/2 Eliminate the cosine from the left hand side. Take the inverse cosine of both sides: x = cos^(-1)(sqrt(37)/2-5/2)+2 pi n_1 for n_1 element Z or x = 2 pi n_2-cos^(-1)(sqrt(37)/2-5/2) for n_2 element Z or 5/2+cos(x) = -sqrt(37)/2 Look at the third equation: Solve for cos(x). Subtract 5/2 from both sides: x = cos^(-1)(sqrt(37)/2-5/2)+2 pi n_1 for n_1 element Z or x = 2 pi n_2-cos^(-1)(sqrt(37)/2-5/2) for n_2 element Z or cos(x) = -5/2-sqrt(37)/2 Show the equation has no solution. cos(x) = -5/2-sqrt(37)/2 has no solution since for all x element C, -1<=Re(cos(x))<=1 and -5/2-sqrt(37)/2<-1: Answer: | | x = cos^(-1)(sqrt(37)/2-5/2)+2 pi n_1 for n_1 element Z or x = 2 pi n_2-cos^(-1)(sqrt(37)/2-5/2) for n_2 element Z